# Online Detecting LLM-Generated Texts via Sequential Hypothesis Testing by Betting

### v2

@@ Appendix B Related Score Functions
The score function $\phi:\text{Text}\rightarrow\mathbb{R}$ will take a text as input and then output a real number. It is designed to maximize the ability to distinguish machine text from human text, that is, we want the score function to maximize the difference in scores between human text and machine text.
@@ Appendix C Proof of Regret Bound of ONS
6:     Get a hint of $d_{t+1}$ to update the domain $\mathcal{K}_{t+1}\leftarrow[-\frac{1}{2d_{t+1}},\frac{1}{2d_{t+1}}]$.
@@
$\theta_{t+1}=\text{proj}_{\mathcal{K}_{t+1}}({\beta}_{t+1})=\arg\min_{\theta\in\mathcal{K}_{t}}(\beta_{t+1}-\theta)^{2}.$
@@
Under $H_{0}$, i.e., $\mu_{x}=\mu_{y}$, the wealth is a P-supermartingale. The value of $g_{t}$ are constrained to the interval $[-1,1]$ in previous works [Orabona and Pál, 2016, Cutkosky and Orabona, 2018, Chugg et al., 2023] with the scores $\phi({x}_{t})$ and $\phi({y}_{t})$ each range from $[0,1]$. To ensure that wealth $W_{t}$ remains nonnegative and to establish the regret bound by ONS, $\theta_{t}$ is always selected within $[-1/2,1/2]$. In our setting, however, the actual range of score difference between two texts, denoted as $g_{t}=\phi({x}_{t})-\phi({y}_{t})$, is typically unknown beforehand. If we assume that the ranges for both $\phi({x}_{t})$ and $\phi({y}_{t})$ are $[m_{t},n_{t}]$, the range of their difference $g_{t}$ is then symmetric about zero, which spans from $-(n_{t}-m_{t})$ to $(n_{t}-m_{t})$. we denote an upper bound $d_{t}\geq n_{t}-m_{t}$ and express the range as $g_{t}\in[-d_{t},d_{t}]$, where $d_{t}\geq 0$. Then, choosing $\theta_{t}$ within $[-1/2d_{t},1/2d_{t}]$ can guarantee that the wealth is a nonnegative P-supermartingale. If we consider the condition of either $W_{t}\geq 1/\alpha\text{ or }W_{T}\geq Z/\alpha$ for any stopping time $T$ as the indication to "reject $H_{0}$" and apply the randomized Ville’s inequality [Ramdas and Manole, 2023], the type-I error can be controlled below the significance level $\alpha$ under $H_{0}$.
Under $H_{1}$, i.e.,$\mu_{x}\neq\mu_{y}$, our goal is to select a proper $\theta_{t}$ at each round $t$ that can speed up the wealth accumulation. It allows us to declare the detection of an LLM once the wealth reaches the specified threshold $1/\alpha$. We can choose $\theta_{t}$ recursively following Algorithm 2. This algorithm can guarantee the regret upper bound for exp-concave loss. Following the proof of Theorem 4.6 in Hazan et al. [2016], we can derive the bound to the regret. The regret of choosing $\theta_{t}\in\mathcal{K}_{t}$ after $T$ time steps by Algorithm 2 is defined as
$\mathrm{Regret}_{T}(\mathrm{ONS}):=\sum_{t=1}^{T}\ell_{t}(\theta_{t})-\sum_{t=1}^{T}\ell_{t}(\theta^{*}),$ (6)
@@
Proof of Lemma 8. The composition of a concave and non-decreasing function with another concave function remains concave. Given that for all $\gamma=\frac{1}{2}\min\left\{\frac{1}{GD},\alpha\right\}$, we have $2\gamma\leq\alpha$, the function $g(\mathbf{\theta})=\mathbf{\theta}^{2\gamma/\alpha}$ composed with $f(\mathbf{\theta})=\exp(-\alpha f(\mathbf{\theta}))$ is concave. Hence, the function $h(\mathbf{\theta})$, defined as $\exp(-2\gamma f(\mathbf{\theta}))$, is also concave. Then by the definition of concavity,
@@
$\exp(-2\gamma f(\mathbf{x}))\leq\exp(-2\gamma f(\mathbf{\beta}))[1-2\gamma\nabla f(\mathbf{\beta})^{\top}(\mathbf{\theta}-\mathbf{\beta})].$ (10)
@@
$f(\mathbf{\theta})\geq f(\mathbf{\beta})-\frac{1}{2\gamma}\log\left(1-2\gamma\nabla f(\mathbf{\beta})^{\top}(\mathbf{\theta}-\mathbf{\beta})\right).$ (11)
Since $D,G$ are previously denoted as the diameter of $\mathcal{K}$ and a bound on the (sub)gradients of $f$ respectively, which means that $D\geq\lvert\mathbf{\theta}-\mathbf{\beta}\rvert$, $G\geq\nabla f(\mathbf{\beta})$. Therefore, we have $|2\gamma\nabla f(\mathbf{\beta})(\mathbf{\theta}-\mathbf{\beta})|\leq 2\gamma GD\leq 1\Rightarrow-1\leq 2\gamma\nabla f(\mathbf{\beta})(\mathbf{\theta}-\mathbf{\beta})|\leq 1$. According to the Taylor approximation, we know that $-\log(1-a)\geq a+\frac{1}{4}a^{2}$ holds for $a\geq-1$. The lemma is derived by considering $a=2\gamma\nabla f(\mathbf{\beta})(\mathbf{\theta}-\mathbf{\beta})$.
@@
We recall that $D_{t}$ is defined as the diameter of $\mathcal{K}_{t}$, i.e., $D_{t}=\max_{a,b\in\mathcal{K}_{t}}\lVert a-b\rVert$, and $G_{t}$ is defined as a bound on the gradients of the loss function $\ell_{t}(\theta)=-\ln(1-g_{t}\theta)$ at time $t$, i.e., $G_{t}=\max_{\theta_{t}\in\mathcal{K}_{t}}\lvert\frac{d}{d\theta_{t}}\ell_{t}(\theta_{t})\rvert$. In our setting, $\mathcal{K}_{t}$ is $[-{1}/{2d_{t}},{1}/{2d_{t}}]$, thus $D_{t}={1}/{d_{t}}$. The gradient $z_{t}=\nabla\ell_{t}(\theta_{t})={g_{t}}/{(1-g_{t}\theta_{t})}$. We find that $\ell_{t}$ monotonically increases with $g_{t}$ and $\theta_{t}$, which means $G_{t}$ can be taken at the maximum $g_{t}$ and the maximum $\theta_{t}$. Since $g_{t}=\phi({x}_{t})-\phi({y}_{t})\in[-d_{t},d_{t}]$, we have $G_{t}={d_{t}}/{(1-d_{t}\cdot\frac{1}{2d_{t}})}=2d_{t}$. Above all, we get $G_{t}D_{t}=2d_{t}\cdot 1/d_{t}=2$ for each $t$, and $\alpha=1$. The value of $\gamma_{t}=\frac{1}{2}\min\{{1}/{G_{t}D_{t}},\alpha\}$ remains a positive constant for all $t$. Therefore, we can simply use $\gamma$ in the following proof since $\gamma_{t}$ is the same for all $t$.
According to the update rule of the algorithm: $\theta_{t+1}=\text{proj}_{\mathcal{K}_{t+1}}(\beta_{t+1})$, and the definition: $\beta_{t+1}=\theta_{t}-\frac{1}{\gamma}\cdot{z_{t}}/{a_{t}}$, we get:
@@
Plugging (17) in (16) gives
$z_{t}(\theta_{t}-\theta^{*})\leq\frac{1}{2\gamma}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}(\theta_{t}-\theta^{*})^{2}a_{t}-\frac{\gamma}{2}(\theta_{t+1}-\theta^{*})^{2}a_{t}.$ (18)
@@
$=\sum_{t=1}^{T}\frac{1}{2\gamma}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}(\theta_{1}-\theta^{*})^{2}a_{1}+\sum_{t=2}^{T}\frac{\gamma}{2}(\theta_{t}-\theta^{*})^{2}(a_{t}-a_{t-1})-\frac{\gamma}{2}(\theta_{T+1}-\theta^{*})^{2}a_{T}\qquad(\text{Since }\gamma>0,a_{T}>0)$
$\leq\sum_{t=1}^{T}\frac{1}{2\gamma}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}(\theta_{1}-\theta^{*})^{2}a_{1}+\sum_{t=2}^{T}\frac{\gamma}{2}(\theta_{t}-\theta^{*})^{2}(a_{t}-a_{t-1})\qquad(\text{Since }a_{t}-a_{t-1}=z_{t}^{2})$
$\leq\sum_{t=1}^{T}\frac{1}{2\gamma}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}(\theta_{1}-\theta^{*})^{2}(a_{1}-z_{1}^{2})+\sum_{t=1}^{T}\frac{\gamma}{2}(\theta_{t}-\theta^{*})^{2}z_{t}^{2}.$
According to the definition: $R_{t}:=z_{t}(\theta_{t}-\theta^{*})-\frac{\gamma}{2}(\theta_{t}-\theta^{*})^{2}z_{t}^{2}$. We move the last term of the right hand side in (LABEL:eq:sum) to the left hand side and get
$\sum_{t=1}^{T}R_{t}\leq\frac{1}{2\gamma}\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}(\theta_{1}-\theta^{*})^{2}(a_{1}-z_{1}^{2}).$ (20)
According to our algorithm, $a_{1}-z_{1}^{2}=a_{0}$. Since $\mathcal{K_{*}}=\bigcap_{t=1}^{T}\mathcal{K}_{t}\subseteq\mathcal{K}_{1}$, the diameter $\|\theta_{1}-\theta^{*}\|^{2}\leq D_{1}^{2}$. We recall the inequality (13), then
$\text{Regret}_{T}(\text{ONS})\leq\sum_{t=1}^{T}R_{t}\leq\frac{1}{2\gamma}\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}D_{1}^{2}a_{0}.$ (21)
@@
$\text{Regret}_{T}(\text{ONS})\leq\frac{1}{2\gamma}\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}D_{1}^{2}a_{0}=\frac{1}{2\gamma}\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}D_{1}^{2}\cdot\frac{1}{\gamma^{2}D_{1}^{2}}=\frac{1}{2\gamma}\left(\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}+1\right).$ (22)
To get the upper bound of regret for our algorithm, we first show that the term $\sum_{t=1}^{T}({z_{t}^{2}}/{a_{t}})$ is upper bounded by a telescoping sum. For real numbers $a,b\in\mathbb{R}_{+}$, the first order Taylor expansion of the natural logarithm of $b$ at $a$ implies ${(a-b)}/{a}\leq\log\left({a}/{b}\right)$, thus
$\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}=\sum_{t=1}^{T}\frac{1}{a_{t}}\cdot(a_{t}-a_{t-1})\leq\sum_{t=1}^{T}\log\left(\frac{a_{t}}{a_{t-1}}\right)=\log\left(\frac{a_{t}}{a_{0}}\right).$ (23)
In our setting, $a_{t}=a_{0}+\sum_{t=1}^{T}z_{t}^{2}$, where $a_{0}=1$, $z_{t}={g_{t}}/{(1-g_{t}\theta_{t})}$. We recall the inequality (21), the upper bound of regret is
$\text{Regret}_{T}(\text{ONS})$ $\leq\frac{1}{2\gamma}\cdot\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}D_{1}^{2}\cdot 1$
$\leq\frac{1}{2\gamma}\cdot\log\left(\frac{a_{t}}{a_{0}}\right)+\frac{\gamma}{2}D_{1}^{2}$
$=\frac{1}{2\gamma}\cdot\log\left(1+\sum_{t=1}^{T}\frac{g_{t}^{2}}{(1-g_{t}\theta_{t})^{2}}\right)+\frac{\gamma}{2}D_{1}^{2}.$ (24)
Since that ${\gamma D_{1}^{2}}$ is a positive constant, $g_{t}\in[-d_{t},d_{t}]$, $\theta_{t}\in[-{1}/{2d_{t}},{1}/{2d_{t}}]$, it follows that $(1-g_{t}\theta_{t})^{2}\in[{1}/{4},{9}/{4}]$. Consequently, we obtain that $\text{Regret}_{T}(\text{ONS})=O\left(\log\left(\sum_{t=1}^{T}g_{t}^{2}\right)^{\frac{1}{2\gamma}}\right)$. This conclusion will be used to show that the update of $\theta_{t}$ in the betting game is to play it on the exp-concave loss $\ell_{t}(\theta)=-\log(1-g_{t}\theta)$ and to get the lower bound of the wealth.
The reason we can obtain the upper bound of regret is that, although the values of $G_{t}$ and $D_{t}$ individually unknown, their product is deterministic. Consequently, the value of $\gamma_{t}=\frac{1}{2}\min\left\{\frac{1}{G_{t}D_{t}},\alpha\right\}$ for all $t$ remains consistent. When we use Lemma 8 to establish the regret bound, as illustrated by equation (LABEL:eq:sum), the uniform $\gamma$ helps us simplify and combine terms to achieve the final result.
Appendix D Lower Bound of the Learner’s Wealth

### v3

@@ Appendix B Related Score Functions
The score function $\phi:\text{Text}\rightarrow\mathbb{R}$ will take a text as input and then output a real number. It is designed to maximize the ability to distinguish machine text from human text, that is, we want the score function to maximize the difference in scores between human-written text and machine-generated text.
@@ Appendix C Proof of Regret Bound of ONS
6:     Get a hint of $d_{t+1}$ to update the decision space $\mathcal{K}_{t+1}\leftarrow[-\frac{1}{2d_{t+1}},\frac{1}{2d_{t+1}}]$.
@@
$\theta_{t+1}=\text{proj}_{\mathcal{K}_{t+1}}({\beta}_{t+1})=\arg\min_{\theta\in\mathcal{K}_{t+1}}(\beta_{t+1}-\theta)^{2}.$
@@
Under $H_{0}$, i.e., $\mu_{x}=\mu_{y}$, the wealth is a P-supermartingale. The value of $g_{t}$ are constrained to the interval $[-1,1]$ in previous works [Orabona and Pál, 2016, Cutkosky and Orabona, 2018, Chugg et al., 2023] with the scores $\phi({x}_{t})$ and $\phi({y}_{t})$ each range from $[0,1]$. To ensure that wealth $W_{t}$ remains nonnegative and to establish the regret bound by ONS, $\theta_{t}$ is always selected within $[-1/2,1/2]$. In our setting, however, the actual range of score difference between two texts, denoted as $g_{t}=\phi({x}_{t})-\phi({y}_{t})$, is typically unknown beforehand. If we assume that the ranges for both $\phi(x_{t})$ and $\phi(y_{t})$ are $[m_{t},n_{t}]$, then the range of their difference $g_{t}:=\phi(y_{t})-\phi(x_{t})$ is symmetric about zero, and is given by the interval $\left[-(n_{t}-m_{t}),;n_{t}-m_{t}\right]$. We denote an upper bound $d_{t}\geq|n_{t}-m_{t}|$ and express the range as $g_{t}\in[-d_{t},d_{t}]$, where $d_{t}\geq 0$. Then, choosing $\theta_{t}$ within $[-1/2d_{t},1/2d_{t}]$ can guarantee that the wealth is a nonnegative P-supermartingale. If we consider the condition of either $W_{t}\geq 1/\alpha\text{ or }W_{T}\geq Z/\alpha$ for any stopping time $T$ as the indication to “reject $H_{0}$" and apply the randomized Ville’s inequality [Ramdas and Manole, 2023], the type-I error can be controlled below the significance level $\alpha$ when $H_{0}$ is true.
Under $H_{1}$, i.e.,$\mu_{x}\neq\mu_{y}$, our goal is to select a proper $\theta_{t}$ at each round $t$ that can speed up the wealth accumulation. It allows us to declare the detection of an LLM once the wealth reaches the specified threshold $1/\alpha$. We can choose $\theta_{t}$ recursively following Algorithm 2. This algorithm can guarantee the regret upper bound for exp-concave loss. Following the proof of Theorem 4.6 in Hazan et al. [2016], we can derive the bound on the regret. The regret of choosing $\theta_{t}\in\mathcal{K}_{t}$ after $T$ time steps by Algorithm 2 is defined as
$\mathrm{Regret}_{T}(\mathrm{ONS}):=\sum_{t=1}^{T}\ell_{t}(\theta_{t})-\sum_{t=1}^{T}\ell_{t}(\theta^{*}),$
@@
Proof of Lemma 5. The composition of a concave and non-decreasing function with another concave function remains concave. Given that for all $\gamma=\frac{1}{2}\min\left\{\frac{1}{GD},\alpha\right\}$, we have $2\gamma\leq\alpha$, the function $g(\mathbf{\theta})=\mathbf{\theta}^{2\gamma/\alpha}$ composed with $f(\mathbf{\theta})=\exp(-\alpha f(\mathbf{\theta}))$ is concave. Hence, the function $h(\mathbf{\theta})=\exp(-2\gamma f(\mathbf{\theta}))$ is also concave. Then by the definition of concavity,
@@
$\exp(-2\gamma f(\theta)\leq\exp(-2\gamma f(\mathbf{\beta}))[1-2\gamma\nabla f(\mathbf{\beta})^{\top}(\mathbf{\theta}-\mathbf{\beta})].$
@@
$f(\mathbf{\theta})\geq f(\mathbf{\beta})-\frac{1}{2\gamma}\ln\left(1-2\gamma\nabla f(\mathbf{\beta})^{\top}(\mathbf{\theta}-\mathbf{\beta})\right).$
Since $D,G$ are previously denoted as the diameter of $\mathcal{K}$ and a bound on the (sub)gradients of $f$ respectively, which means that $D\geq\lvert\mathbf{\theta}-\mathbf{\beta}\rvert$, $G\geq\nabla f(\mathbf{\beta})$. Therefore, we have $|2\gamma\nabla f(\mathbf{\beta})(\mathbf{\theta}-\mathbf{\beta})|\leq 2\gamma GD\leq 1\Rightarrow-1\leq 2\gamma\nabla f(\mathbf{\beta})(\mathbf{\theta}-\mathbf{\beta})\leq 1$. According to the Taylor approximation, we know that $-\ln(1-a)\geq a+\frac{1}{4}a^{2}$ holds for $a\geq-1$. The lemma is derived by considering $a=2\gamma\nabla f(\mathbf{\beta})(\mathbf{\theta}-\mathbf{\beta})$.
@@
We recall that $D_{t}$ is defined as the diameter of $\mathcal{K}_{t}$, i.e., $D_{t}=\max_{a,b\in\mathcal{K}_{t}}|a-b|$, and $G_{t}$ is defined as a bound on the gradients of the loss function $\ell_{t}(\theta)=-\ln(1-g_{t}\theta)$ at time $t$, i.e., $G_{t}=\max_{\theta_{t}\in\mathcal{K}_{t}}\lvert\frac{d}{d\theta_{t}}\ell_{t}(\theta_{t})\rvert$. In our setting, $\mathcal{K}_{t}$ is $[-{1}/{2d_{t}},{1}/{2d_{t}}]$, thus $D_{t}={1}/{d_{t}}$. The gradient $z_{t}=\nabla\ell_{t}(\theta_{t})={g_{t}}/{(1-g_{t}\theta_{t})}$, so we have $G_{t}=\max_{\theta_{t}\in\left[-\frac{1}{2d_{t}},\frac{1}{2d_{t}}\right]}\frac{|g_{t}|}{1-g_{t}\theta_{t}}$. Since $g_{t}=\phi({x}_{t})-\phi({y}_{t})\in[-d_{t},d_{t}]$, and the expression $\frac{|g_{t}|}{1-g_{t}\theta_{t}}$ increases with the product $g_{t}\theta_{t}\in[-1/2,1/2]$, the maximum is attained when $g_{t}=d_{t}$ and $\theta_{t}=\frac{1}{2d_{t}}$. Thus, we obtain $G_{t}={d_{t}}/{(1-d_{t}\cdot\frac{1}{2d_{t}})}=2d_{t}$. Above all, we get $G_{t}D_{t}=2d_{t}\cdot\frac{1}{d_{t}}=2$ for each $t$, and $\alpha=1$. The value of $\gamma_{t}=\frac{1}{2}\min\{{1}/{G_{t}D_{t}},\alpha\}$ remains a positive constant for all $t$. Therefore, we simply use $\gamma$ in the following proof, given that $\gamma_{t}$ takes the same constant value for all $t$.
According to the update rule of the algorithm: $\theta_{t+1}=\text{proj}_{\mathcal{K}_{t+1}}(\beta_{t+1})$, and the definition: $\beta_{t+1}=\theta_{t}-\frac{1}{\gamma}\cdot\frac{z_{t}}{a_{t}}$, we get:
@@
Plugging (12) into (11) and rearranging terms yields
$z_{t}(\theta_{t}-\theta^{*})\leq\frac{1}{2\gamma}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}(\theta_{t}-\theta^{*})^{2}a_{t}-\frac{\gamma}{2}(\theta_{t+1}-\theta^{*})^{2}a_{t}.$
@@
$=\sum_{t=1}^{T}\frac{1}{2\gamma}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}(\theta_{1}-\theta^{*})^{2}a_{1}+\sum_{t=2}^{T}\frac{\gamma}{2}(\theta_{t}-\theta^{*})^{2}(a_{t}-a_{t-1})-\frac{\gamma}{2}(\theta_{T+1}-\theta^{*})^{2}a_{T}$
$\leq\sum_{t=1}^{T}\frac{1}{2\gamma}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}(\theta_{1}-\theta^{*})^{2}a_{1}+\sum_{t=2}^{T}\frac{\gamma}{2}(\theta_{t}-\theta^{*})^{2}(a_{t}-a_{t-1})\qquad(\text{since }\gamma>0,a_{T}>0)$
$\leq\sum_{t=1}^{T}\frac{1}{2\gamma}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}(\theta_{1}-\theta^{*})^{2}(a_{1}-z_{1}^{2})+\sum_{t=1}^{T}\frac{\gamma}{2}(\theta_{t}-\theta^{*})^{2}z_{t}^{2}\qquad(\text{by }a_{t}-a_{t-1}=z_{t}^{2}).$ (13)
According to the definition: $R_{t}:=z_{t}(\theta_{t}-\theta^{*})-\frac{\gamma}{2}(\theta_{t}-\theta^{*})^{2}z_{t}^{2}$. We move the last term of the right hand side in (13) to the left hand side and get
$\sum_{t=1}^{T}R_{t}\leq\frac{1}{2\gamma}\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}(\theta_{1}-\theta^{*})^{2}(a_{1}-z_{1}^{2}).$
According to our algorithm, $a_{1}-z_{1}^{2}=a_{0}$. Since $\mathcal{K_{*}}=\bigcap_{t=1}^{T}\mathcal{K}_{t}\subseteq\mathcal{K}_{1}$, the diameter $(\theta_{1}-\theta^{*})^{2}\leq D_{1}^{2}$. We recall the inequality (8), then
$\text{Regret}_{T}(\text{ONS})\leq\sum_{t=1}^{T}R_{t}\leq\frac{1}{2\gamma}\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}(\theta_{1}-\theta^{*})^{2}a_{0}\leq\frac{1}{2\gamma}\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}D_{1}^{2}a_{0}.$ (14)
@@
$\text{Regret}_{T}(\text{ONS})\leq\frac{1}{2\gamma}\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}D_{1}^{2}a_{0}=\frac{1}{2\gamma}\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}D_{1}^{2}\cdot\frac{1}{\gamma^{2}D_{1}^{2}}=\frac{1}{2\gamma}\left(\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}+1\right).$
To get the upper bound of regret for our algorithm, we first show that the term $\sum_{t=1}^{T}({z_{t}^{2}}/{a_{t}})$ is upper bounded by a telescoping sum. For real numbers $a,b\in\mathbb{R}_{+}$, the first order Taylor expansion of the natural logarithm of $b$ at $a$ implies $\frac{a-b}{a}\leq\ln\frac{a}{b}$, thus
$\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}=\sum_{t=1}^{T}\frac{1}{a_{t}}\cdot(a_{t}-a_{t-1})\leq\sum_{t=1}^{T}\ln\left(\frac{a_{t}}{a_{t-1}}\right)=\ln\left(\frac{a_{T}}{a_{0}}\right).$
In our setting, $a_{T}=a_{0}+\sum_{t=1}^{T}z_{t}^{2}$, where $a_{0}=1$, $z_{t}=\frac{g_{t}}{1-g_{t}\theta_{t}}$. We recall the inequality (14), the upper bound of regret is
$\text{Regret}_{T}(\text{ONS})$ $\leq\frac{1}{2\gamma}\cdot\sum_{t=1}^{T}\frac{z_{t}^{2}}{a_{t}}+\frac{\gamma}{2}(\theta_{1}-\theta^{*})^{2}\cdot 1$
$\leq\frac{1}{2\gamma}\cdot\ln\left(\frac{a_{T}}{a_{0}}\right)+\frac{\gamma}{2}D_{1}^{2}$
$=\frac{1}{2\gamma}\cdot\ln\left(1+\sum_{t=1}^{T}\frac{g_{t}^{2}}{(1-g_{t}\theta_{t})^{2}}\right)+\frac{\gamma}{2}D_{1}^{2}$
$\leq\frac{1}{2\gamma}\cdot\ln\left(1+4\sum_{t=1}^{T}g_{t}^{2}\right)+\frac{\gamma}{2}D_{1}^{2}$
$\leq\frac{1}{2\gamma}\cdot\ln\left(1+4d_{*}^{2}T\right)+\frac{\gamma}{2}D_{1}^{2}.$ (15)
Since that ${\gamma D_{1}^{2}}$ is a positive constant, $g_{t}\in[-d_{t},d_{t}]$, $\theta_{t}\in[-{1}/{2d_{t}},{1}/{2d_{t}}]$, it follows that $(1-g_{t}\theta_{t})^{2}\in[\frac{1}{4},\frac{9}{4}]$. We denote $d_{*}:=\max_{t\geq 1}|d_{t}|$, which indicates that $g_{t}^{2}\leq d_{t}^{2}\leq d_{*}^{2}$. Consequently, we obtain that $\text{Regret}_{T}(\text{ONS})=\mathcal{O}\left(\ln T^{\frac{1}{2\gamma}}\right)$. This conclusion will be used to show that the update of $\theta_{t}$ in the betting game is to play it against the exp-concave loss $\ell_{t}(\theta)=-\ln(1-g_{t}\theta)$ and to get the lower bound of the wealth.
The reason we can obtain the upper bound of regret is that, although the values of $G_{t}$ and $D_{t}$ individually unknown, their product is deterministic. Consequently, the value of $\gamma_{t}=\frac{1}{2}\min\left\{\frac{1}{G_{t}D_{t}},\alpha\right\}$ for all $t$ remains consistent. When we use Lemma 5 to establish the regret bound, as illustrated by equation (13), the uniform $\gamma$ helps us simplify and combine terms to achieve the final result.
Appendix D Lower Bound on the Learner’s Log-Wealth